# Adaptive Hyperparameter Optimization for Continual Learning Scenarios

### v1

@@ 2.3.2 Related Works
Hyperparameters play a significant role in continual learning by influencing the model’s ability to adapt to new tasks while retaining knowledge from previous tasks. Note that strategies tackling the continual learning problem typically involve extra hyperparameters to balance the stability plasticity trade-off. These hyperparameters are in many cases found via a grid search, using held-out validation data from all tasks [6]. In particular, the typical assumption is to have access to the entire stream at the end of training for model selection purposes. After that, as a common practice, the hyperparameters are set to a fixed value throughout all incremental learning sessions. One simple motivation is that tuning hyperparameters is a major burden and is no different in the continual setting. Another motivation is the simplicity of implementation. Nevertheless, this inherently violates the main assumption in continual learning, namely no access to previous task data and in practical scenarios is unrealistic for building accurate lifelong learning machines [10, 5].
In the following, we present a comprehensive summary of related works concerning adaptive HPO and CL, categorizing them into three classes.
@@ 3.1 Methodology and Key Assumptions
[…]
@@
8:     elsenormal-$\triangleright$ Rest of other tasks
@@
hpo: Implement a specific HPO for each model update, i.e. grid search, population-based training or Bayesian optimization. The latter should speed up the HPO if we start from the best configuration in the previous task. For this reason, we want to use Tree-structured Parzen Estimator (TPE) [2] that performs robust and efficient hyperparameter optimization at scale by combining the speed of Hyperband searches with the guidance and guarantees of convergence of Bayesian Optimization among tasks.
@@ 3.2.1 Benchmarks and Metric
In this paper, we would conduct experiments using benchmarks from two distinct scenarios: Class Incremental, where each task introduces new classes without revisiting old ones in the training stream, and Domain Incremental, where each task presents new instances for existing classes without reusing old instances in the training stream [15]. We have selected three for Class Incremental Learning: Split-CIFAR10 [31], Split-TinyImageNet [21] and CORe50-NC [18]. These benchmarks are derived respectively from CIFAR-10 [13] and TinyImagenet [14] datasets while CORe50-NC is a benchmark specifically designed for continual learning. For Domain Incremental Learning we have selected Rotated-MINIST derived from the MNIST dataset [7] and CORe50-NI.
Each dataset exhibits visual classification learning. We will train all the models with Split-CIFAR10 and Rotated-MINIST online, while the others will be in batch mode. We will resort to the standard metrics for evaluation, i.e. accuracy, which measures the final performance averaged over all tasks (also defined as stream accuracy, SA) [15]. In both incremental scenarios, higher values indicate better performance.
@@ 3.2.2 Baselines
  • The upper bound in terms of performance and lower bound for computational cost is HPO for each model update as used in [30]. We plan to use grid search or Bayesian optimization as an HPO technique to improve as possible the performance comparison. The computational cost expected to achieve is ||*$|\mathcal{H}|*|\mathcal{T}|$ which is the worst case.
@@ 5 Changes and Additions to the pre-registered proposal
Given the time constraints and our obtained results, we have chosen to conduct a comprehensive analysis of hybrid and traditional continual learning strategies, emphasizing their robustness in per-task performance in conjunction with HPO. This decision involved focusing on four benchmarks, i.e. Rot-MNIST, Split-CIFAR10, Split-Tiny and CORe50 within a DIL setting. Moreover, the effort made in exploring the importance of hyperparameters for these four benchmarks has improved the analysis of how to tune the parameters better for hybrid and traditional continual learning strategies.
We have also adjusted the definition of our two baselines and their comparison role concerning our method.
  • Fixed. The first baseline random choice of the hyperparameters in the first task and kept fixed in the sequence. We consider this baseline the lower bound in terms of performance and call it Fixed for the rest of the paper.
  • HPO. This second baseline performs the tuning of hyperparameters per task using TPE as an optimization strategy. This baseline is the upper bound in terms of accuracy but with a high computational cost, that we aim to reduce with our approach. We call it HPO for the rest of the paper.

### v2

@@ 2.3.2 Related Works
Hyperparameters play a significant role in continual learning by influencing the model’s ability to adapt to new tasks while retaining knowledge from previous tasks. Note that strategies tackling the continual learning problem typically involve extra hyperparameters to balance the stability plasticity trade-off. These hyperparameters are in many cases found via a grid search, using held-out validation data from all tasks [6]. In particular, the typical assumption is to have access to the entire stream at the end of training for model selection purposes. After that, as a common practice, the hyperparameters are set to a fixed value throughout all incremental learning sessions. One simple motivation is that tuning hyperparameters is a major burden and is no different in the continual setting. Another motivation is the simplicity of implementation. Nevertheless, this inherently violates the main assumption in continual learning, namely no access to previous task data and in practical scenarios is unrealistic for building accurate lifelong learning machines [11, 5]. In the following, we present a comprehensive summary of related works concerning adaptive HPO and CL, categorizing them into three classes.
@@ 3.1 Methodology and Key Assumptions
6:         $\{\mathcal{H}_{n},\mathrm{v}\}=\text{get\_param\_imp}(\text{\emph{f}ANOVA},\text{hpo},\mathcal{H})$
@@
8:     $\triangleright$ Rest of other tasks
@@
hpo: Implement a specific HPO for each model update, i.e. grid search, population-based training or Bayesian optimization. The latter should speed up the HPO if we start from the best configuration in the previous task. For this reason, we want to use Bayesian Optimization and Hyperband (BOHB) [8] that performs robust and efficient hyperparameter optimization at scale by combining the speed of Hyperband searches with the guidance and guarantees of convergence of Bayesian Optimization among tasks.
@@ 3.2.1 Benchmarks and Metric
In this paper, we would conduct experiments using benchmarks from two distinct scenarios: Class Incremental, where each task introduces new classes without revisiting old ones in the training stream, and Domain Incremental, where each task presents new instances for existing classes without reusing old instances in the training stream [16]. We have selected three for Class Incremental Learning: Split-CIFAR10 [32], Split-TinyImageNet [22] and CORe50-NC [19]. These benchmarks are derived respectively from CIFAR-10 [14] and TinyImagenet [15] datasets while CORe50-NC is a benchmark specifically designed for continual learning. For Domain Incremental Learning we have selected Rotated-MINIST derived from the MNIST dataset [7] and CORe50-NI. Each dataset exhibits visual classification learning. We will train all the models with Split-CIFAR10 and Rotated-MINIST online, while the others will be in batch mode. We will resort to the standard metrics for evaluation, i.e. accuracy, which measures the final performance averaged over all tasks (also defined as stream accuracy, SA) [16]. In both incremental scenarios, higher values indicate better performance.
@@ 3.2.2 Baselines
  • The upper bound in terms of performance and lower bound for computational cost is HPO for each model update as used in [31]. We plan to use grid search or Bayesian optimization as an HPO technique to improve as possible the performance comparison. The computational cost expected to achieve is $|\mathcal{H}|*|\mathcal{T}|$ which is the worst case.
@@ 5 Changes and Additions to the pre-registered proposal
Given the time constraints and our obtained results, we have chosen to conduct a comprehensive analysis of hybrid and traditional continual learning strategies, emphasizing their robustness in per-task performance in conjunction with HPO. This decision involved focusing on four benchmarks, i.e. Rot-MNIST, Split-CIFAR10, Split-Tiny and CORe50 within a DIL setting. Moreover, the effort made in exploring the importance of hyperparameters for these four benchmarks has improved the analysis of how to tune the parameters better for hybrid and traditional continual learning strategies. Moreover, we have decided to change the optimizer, specifically the Tree-structured Parzen Estimator (TPE) [2], instead of BOHB. The main reason for this change is that TPE is highly efficient in searching for hyperparameters and can find good sets with relatively few evaluations compared to BOHB. Finally, we have named the two baselines. The first baseline, which represents the upper bound in terms of accuracy but comes with a high computational cost, is called HPO while the second baseline is called Fixed.